A new furnace for your small factory will cost $27,000 to install and will require ongoing maintenance expenditures of $1,500 a year. But it is far more fuelefficient than your old furnace and will reduce your consumption of heating oil by 2,400 gallons per year. Heating oil this year will cost $3 a gallon; the price per gallon is expected to increase by $.50 a year for the next 3 years and then to stabilize for the foreseeable future. The furnace will last for 20 years, at which point it will need to be replaced and will have no salvage value. The discount rate is 8%. (What is the equivalent annual cost of the furnace?) What is the equivalent annual savings derived from the furnace?
Answer #1

EAC=(1500+27000)*8%/(1-1/1.08^20)=2902.787951

EAS=(2400*3/1.08+2400*3.5/1.08^2++2400*4/1.08^3+2400*4.5/1.08^4+2400*5/1.08^5*(1-1/1.08^16))*8%/(1-1/1.08^20)=3586.272706